In Africa in 2011, the number of deaths of a female from cardiovascular disease for different age groups are in table #11.2.6 ("Global health observatory," 2013). In addition, the proportion of deaths of females from all causes for the same age groups are also in table #11.2.6. Do the data show that the death from cardiovascular disease are in the same proportion as all deaths for the different age groups? Test at the 5% level.

Table #11.2.6: Deaths of Females for Different Age Groups

Age

5-14

15-29

30-49

50-69

Total

Cardiovascular Frequency

8

16

56

433

513

All Cause Proportion

0.10

0.12

0.26

0.52
Answer #1

Solution: H.:P5-14 = 0.10, P15-29 = 0.12, P30-49 = 0.26, P50-69 = 0.52 Hj: The proportions are not p5-14 = 0.10, P15-29 = 0.1Since the value of test statistic x2 = 218.756 falls in the rejection region, we reject the null hypothesis. Conclusion: Reje